# The Learning Approach to Games

### v1

@@ 2.1 Two player game example
Consider a fixed state and action space as $\mathbb{S}=\{0,1\}$ and $\mathbb{A}=[0,1]$. Suppose players are not learning the horizon and transition probabilities, that is, $\mathfrak{L}_{T}^{i},\mathfrak{L}_{p}^{i}$ are constants yielding $\hat{T}^{i}=1$ and $p^{i}(t,\vec{x},\vec{a},1)=a^{i}$. Initially, for an easier comparison with the Nash equilibrium, we assume player’s are not learning the state value and transition costs either, and it is given by
@@ 2.1.1 Details of the Learning Algorithm
We now introduce the relevant parts of the framework specific for this problem. In general, each player models a transition cost $\hat{F}$ and an estimate $\hat{\Gamma}$ on the other player’s actions, relying on their observations held in the memory.
@@
There are two objectives cost networks are training for; (i) there is an expected cost coming from the predictions of action networks, which is (2.15) integrated as in (2.6). If action networks are not perfect, expected cost will not match the observed expected cost. Cost networks are training to close this gap, by relying on a cost memory similar to action memory. (ii): players have expectations over what is best possible as introduced in (2.13), which we took as constant for simplicity. Cost networks are also training such that the players do not get desperate. In the simulations, if the first player ends up with networks ${\cal N}_{\Gamma}^{1,k}$’s taking values close to $1$, independent of the action, they start to play the Nash equilibrium $(1,1)$. That is, player $1$ gets desperate, and then adjusts the random component of cost to start explore other actions.
@@
Before discussing the simulation results, let us mention that each parameter of networks of course plays a significant role, and we coarsely tuned them by hand to obtain simulations matching with expectations. Many of such parameters are taken as constant. Changing to different constants might of course yield poor results. On the other hand, generalizing them will improve the sophistication of agents strategies. Besides the network parameters, there are more structural parameters too. For example, what players are expecting as the best possible cost also changes the characteristics of actions. Especially if the first player is expecting much better than what is realistically possible, exploration gets out of hand and they don’t converge anywhere. For an another example, if the memory is very long and not forgetting, then the first player starts to get advantage over the second player, as the second player becomes fixed after a while and ”thinks” that the other player will still frequently move to state 00. The point we are aiming to convey here is that the learning algorithms of players are crucial to characterize what is going to be realized, and the evolution of ${}^{n}\gamma_{\mathfrak{L}}^{i}\in{\cal P}([0,1])$ might be dynamic even in the simplest settings.
@@
We would like to conclude this section by emphasizing again that games are inherently complex and that observed behaviors require a more detailed representation of players. We refer to the supplementary online repository [6] for the animation showing the cost networks, action networks, and other observables in each case.
3 Stated Mean Field Games
In this section, we will introduce a mean field type version of the framework. It is important to note that learning parameters are defined for each player individually. Therefore, embedding a mean field game requires adjusting the learning parameters of a representative agent to model infinitely many similar players. In particular, to align with a similar structure in the literature, we assume that essentially only $\hat{\Gamma}$ will be learned, while other parameters $(\hat{T},\hat{p},\hat{F},\hat{\phi})=(T,p,F,\phi)$ are modeled by the representative player as known (and not learned). We will thus refer to this case as the stated mean field game.
@@ 3 Stated Mean Field Games
Next, along the idea that the representative agent is insignificant in the population, we assume $\hat{\Gamma}$ in (2.3) is constant as $\hat{\Gamma}:\mathbb{T}\to{\cal P}({\cal P}(\mathbb{S}_{t}\times{\cal A}))$. Here, ${\cal P}(\mathbb{S}_{t}\times{\cal A})$ corresponds to $\vec{\cal A}$ in (2.3). In the case of countable players, indexing was keeping track of the connection between the state and the control of individiual players. Here, state variable keeps track of distribution of controls used by the population.
@@
Finally, we are ready to introduce the crucial part: observations and the learning algorithm. Let us briefly recap the mean field framework. We assume that the representative player starts with an initial guess of the population distribution over states and controls, determines the corresponding optimal control, and, relying on the assumption that everyone else is exactly the same, generates further observations using the chosen learning algorithm. To formalize this, let ${\cal E}={\cal P}({\cal P}(\mathbb{S}\times{\cal A}))$ be the space of observables, and let $\mathscr{E}$ denote the space of finite sequences of ${\cal E}$. Recall that the learning algorithm is a mapping $\mathfrak{L}_{\varphi}:\mathscr{E}\to\mathscr{M}_{\varphi}$, where the player’s estimation at age $n$ is ${}^{n}\mathfrak{L}_{\varphi}:=\mathfrak{L}_{\varphi}({}^{n}{\cal O})$, and the player’s prior is ${}^{0}\mathfrak{L}_{\varphi}$. Here, ${\cal O}:\Omega^{u}\times\mathbb{N}\to\mathscr{E}$ represents the increasing sequence of observations. The set $\Omega^{u}$ corresponds to a potential random choice made by the player during the search for optimal control. To provide intuition, we will construct a simple but explicit learning algorithm for $\mathfrak{L}_{\Gamma}$:
@@
Let us remark on the similarity between the fictitious play-type algorithms introduced in [4]. In fictitious play, one also starts with an initial guess $\delta_{{}^{0}\alpha}$ and finds the optimal ${}^{1}\alpha$. A crucial difference, however, is that fictitious play considers the weighted average of $\mu^{{}^{0}\Xi}$ and $\mu^{{}^{1}\Xi}$ to find the next optimal control $\alpha^{2}$. That is, the cost structure becomes
@@
Definition 4 (Uncertain Equilibria of stated Mean Field Games)
@@ 4 Reinforcement Learning
In this section, we consider the special case of the framework presented in Section 2 when applied to a single player, thereby reducing the problem to a control setting. The main objective is to demonstrate that our framework generalizes Markov Decision Processes, and consequently, well-established methods from the reinforcement learning (RL) literature can be naturally incorporated.
Let us first note that the standard RL terminologies correspond to the learning parameters we have introduced. Learning $\hat{p}$ is typically carried out using model-based methods. In its simplest form, learning the horizon $\hat{T}$ is achieved by adjusting the discount factor [^10] We note that considering an infinite time horizon with exponentially decaying bounded rewards is essentially equivalent to a finite time problem without considering state values.; in more advanced frameworks, it is modeled as a stopping time and incorporated within the options framework. The parameter $\hat{F}$, which represents the rewards, may either be predefined or learned through various methods, such as inverse reinforcement learning. The function $\hat{\phi}$ plays a similar role to the standard value function and is considered distributional RL because it is treated as a random variable. Learning $\hat{\pi}$ is typically referred to as policy learning, while $\hat{\Gamma}$ is sometimes called behavior prediction. It is important to note that an agent may employ more sophisticated strategies by modeling additional, related learning parameters. For example, this could include learning to encode raw observations, establishing communication protocols, understanding opponents’ incentives, or tuning exploration parameters, such as the best expected rewards in the two-player game example in Section 2.1.
@@
For notation, we didn’t keep track of learning index $n$, but both $\tilde{\mathbb{P}}$ and $\hat{F}$ (and consequently $Q$) are evolving as we learn. Note that $Q$ satisfies a Bellman type relation
@@
assigns a scalar value for $(x,a)$, which is better suited in our framework as it does not introduce simplifications. We also note that, in simpler settings, it might be more appropriate to learn an estimate $Q(x,a)$ directly, and avoid all the estimations we have introduced. This is of course a design choice for the player.
We will briefly illustrate a simple method for learning CartPole [7], aiming to further highlight the connections with well-known methods, and learn $\hat{\phi}$ with a fixed horizon $T=8$. Let us revisit the basics of the CartPole problem. The state space is $\mathbb{S}=\mathbb{R}^{4}$ representing position, velocity, angle, and angular velocity. The action space is $\mathbb{A}=\{0,1\}$, where the actions represent applying force to the left or right of the cart. The player gains a reward at each time step if the pole is vertical within some bounds.
There is a memory of player for recording observations: states, actions and a metric evaluating the performance of the player after each episode. We took this performance metric as an average of both relative and absolute performance.
@@
To make sure that we are training value networks properly  [^11] Note that, it is considerably easier to learn controls in this simple setting since $\mathbb{A}^{T}$ is finite, and we observed the player performs quite well even without properly trained value networks., we consider the set of all controls $\mathbb{A}^{T}=\{0,1\}^{T}$ and index them by $\{\alpha^{k}\}_{k=1}^{2^{T}}$. Then, we let $\mathfrak{L}_{\pi}:\mathscr{E}\to(\hat{\Omega}\times\mathbb{S}\to{\cal P}(\mathbb{A}^{T}))$ as
@@
Lastly, we will briefly examine the simplest case of Bandit Games and demonstrate how it can be embedded into our framework in two different ways. The first approach adopts a single-player perspective, where the arms are represented by the state $X$. The second approach treats arms as players who take actions but do not engage in learning. Both versions will help to explore our framework further, and they may generalize in different directions.
@@
Observe that, if the expected reward of one arm is high enough such that the randomness of $\hat{F}$ has support smaller than the difference between their expected rewards, then the optimal induced action becomes a Dirac distribution. However, typically (2.9) induces a non-degenerate distribution over actions, as it is not known by the agent which arm is definitively better. Note that determining the optimal action per event $\hat{\omega}$ is trivial, hence we do not discuss $\hat{\pi}$. Additionally, $\hat{F}$ might aim to capture some related rewards rather than the underlying distribution. For example, in the case of two arms with identical expected rewards, the agent might aim to capture different qualitative differences between the distributions.
@@
In the action perspective, let players $1,2,\dots,k$ represent arms in the game. Each arm has a fixed horizon $\hat{T}=1$, and transition probabilities or estimating others’ actions are not relevant for the arms. We characterize the reward of the main agent as the action taken by the chosen arm. For the rewards of the arms, one can define $\hat{F}^{i}(\hat{\omega},a)$ on a probability space $(\hat{\Omega},\hat{\mathbb{P}})$ such that the induced distributions over actions correspond to the targeted reward distribution $\mu^{i}$. Note that, as the arms do not engage in learning (constant $\mathfrak{L}_{\varphi}$), they always satisfy the conditions of uncertain equilibrium.

### v2

@@ 2.1 Two player game example
Consider a fixed state and action space as $\mathbb{S}=\{0,1\}$ and $\mathbb{A}=[0,1]$. Suppose players are not learning the horizon and transition probabilities, that is, $\mathfrak{L}_{T}^{i},\mathfrak{L}_{p}^{i}$ are constants yielding $\hat{T}^{i}=1$ and $p^{i}(t,\vec{x},\vec{a},1)=a^{i}$. Initially, to state the rules of the game and to enable comparison with the Nash equilibrium, we assume players are not learning the state value or transition costs;
@@ 2.1.1 Details of the Learning Algorithm
We now introduce the relevant parts of the framework specific for this problem. In general, each player models a transition cost $\hat{F}$ and an estimate $\hat{\Gamma}$ on the other player’s actions, relying on their observations held in the memory. Also, each player has an expectation defined as in (2.13).
@@
There are two objectives cost networks are training for: (i) there is an expected cost coming from the predictions of action networks, which is (2.15) integrated as in (2.6). If action networks are not perfect, expected cost will not match the observed expected cost. Cost networks are training to close this gap, by relying on a cost memory similar to action memory. (ii): players have expectations over what is best possible as introduced in (2.13), which we took as constant for simplicity. Cost networks are also trained such that the players do not get desperate. In the simulations, if the first player ends up with networks ${\cal N}_{\Gamma}^{1,k}$’s taking values close to $1$, independent of the action, they start to play the Nash equilibrium $(1,1)$. That is, player $1$ gets desperate, and then adjusts the random component of cost to start exploring other actions.
@@
Before discussing the simulation results, let us mention that each parameter of networks of course plays a significant role, and we coarsely tuned them by hand to obtain simulations matching with expectations. Many of such parameters are taken as constant. Changing to different constants might of course yield poor results. On the other hand, generalizing them will improve the sophistication of agents strategies. Besides the network parameters, there are more structural parameters too. For example, what players are expecting as the best possible cost also changes the characteristics of actions. Especially if the first player is expecting much better than what is realistically possible, exploration gets out of hand and they don’t converge anywhere. For an another example, if the memory is very long and not forgetting, then the first player starts to get advantage over the second player, as the second player becomes fixed after a while and ”thinks” that the other player will still frequently move to state 00. The point we are aiming to convey here is that the learning algorithms of players are crucial to characterize what is going to be realized in reality, and the evolution of ${}^{n}\gamma_{\mathfrak{L}}^{i}\in{\cal P}([0,1])$ might be dynamic even in the simplest settings.
@@
We refer to the supplementary online repository [20] for the animation showing the cost networks, action networks, and other observables in each case.
We conclude this section by emphasizing once more that games are inherently complex and that observed behaviors require a detailed representation of players. In our working paper [4], we explore firm collusion in an online price competition setting. In Calvano et al. [8], the authors model each firm using $Q$-learning. Each firm maintains a single $Q$-function estimate, requiring an intensive training period to populate the $Q$ table, an approach that is infeasible in online learning situations and offers little intuition about why collusion emerges. In [4], we design firms as outlined in this section. Beyond observing rapid convergence, we aim to explain how collusion, or competition, is driven by firms’ design choices, potentially offering policymakers better guidance for regulating price competition.
3 Stated Mean-Field Games
In this section, we will introduce a mean-field type version of the framework. It is important to note that learning parameters are defined for each player individually. Therefore, embedding a mean-field game requires adjusting the learning parameters of a representative agent to model infinitely many similar players. In particular, to align with a similar structure in the literature, we assume that only $\hat{\Gamma}$ will be learned, while other parameters $(\hat{T},\hat{p},\hat{F},\hat{\phi})=(T,p,F,\phi)$ are modeled by the representative player as known (and not learned). We will thus refer to this case as the stated mean-field game.
@@ 3 Stated Mean-Field Games
Next, along the idea that the representative agent is insignificant in the population, we assume $\hat{\Gamma}$ in (2.4) is constant as $\hat{\Gamma}:\mathbb{T}\to{\cal P}({\cal P}(\mathbb{S}_{t}\times{\cal A}))$. Here, ${\cal P}(\mathbb{S}_{t}\times{\cal A})$ corresponds to $\vec{\cal A}$ in (2.4). In the case of countable players, indexing was keeping track of the connection between the state and the control of individual players. Here, state variable keeps track of distribution of controls used by the population.
@@
Finally, we are ready to introduce the crucial part: observations and the learning algorithm. Let us briefly recap the mean-field framework. We assume that the representative player starts with an initial guess of the population distribution over states and controls, determines the corresponding optimal control, and, relying on the assumption that everyone else is exactly the same, generates further observations using the chosen learning algorithm. To formalize this, let ${\cal E}={\cal P}({\cal P}(\mathbb{S}\times{\cal A}))$ be the space of observables, and let $\mathscr{E}$ denote the space of finite sequences of ${\cal E}$. Recall that the learning algorithm is a mapping $\mathfrak{L}_{\varphi}:\mathscr{E}\to\mathscr{M}_{\varphi}$, where the player’s estimation at age $n$ is ${}^{n}\mathfrak{L}_{\varphi}:=\mathfrak{L}_{\varphi}({}^{n}{\cal O})$, and the player’s prior is ${}^{0}\mathfrak{L}_{\varphi}$. Here, ${\cal O}:\Omega^{u}\times\mathbb{N}\to\mathscr{E}$ represents the increasing sequence of observations. The set $\Omega^{u}$ corresponds to a potential random choice made by the player during the search for optimal control. To provide intuition, we will construct a simple but explicit learning algorithm for $\mathfrak{L}_{\Gamma}$:
@@
Let us remark on the similarity between the fictitious play-type algorithms introduced in Cardaliaguet-Hadikhanloo [11]. In fictitious play, one also starts with an initial guess $\delta_{{}^{0}\alpha}$ and finds the optimal ${}^{1}\alpha$. A crucial difference, however, is that fictitious play considers the weighted average of $\mu^{{}^{0}\Xi}$ and $\mu^{{}^{1}\Xi}$ to find the next optimal control $\alpha^{2}$. That is, the cost structure becomes
@@
Definition 4 (Uncertain Equilibria of stated Mean-Field Games)
@@ 4 Reinforcement Learning
We now specialize the framework of Section 2 to a single player, reducing the problem to a control setting. The main objective is to show that familiar reinforcement learning (RL) methods naturally fall within the scope of this framework. Many RL designs can be viewed as instantiations of the player defined in Definition 1.  [^11] In evolutionary approaches, it is more natural to work with a population of players, as in Section 2. However, the evolutionary mechanisms needed are not formalized in this work.
The simplest illustration is vanilla $Q$-learning. The agent maintains a single space of estimations
$\mathscr{M}_{Q}:=\big{\{}\,Q:\mathbb{S}\times\mathbb{A}\to\mathbb{R}\,\big{\}}$
and, after training, adjusts the behavior to act greedily:
${\Upsilon}:\mathscr{M}_{Q}\to{\cal P}(\mathbb{A})\qquad{\Upsilon}(Q)=\delta_{\operatorname*{arg\,max}_{a\in\mathbb{A}}Q(x,a)}$
The learning algorithm $\mathfrak{L}\colon\mathscr{E}\to\mathscr{M}_{Q}$ implemented, for example, via temporal-difference or Monte-Carlo updates, iteratively adjusts the estimation to yield higher values for favorable outcomes, while preserving the time-consistency of the estimation.
We note that the standard RL terminology maps naturally to the learning parameters we have introduced. Learning $\hat{p}$ is typically associated with model-based methods. The horizon $\hat{T}$ is often handled by adjusting the discount factor, though more advanced approaches model it as a stopping time within the options framework. The parameter $\hat{F}$, representing rewards, may be predefined or learned through methods such as inverse reinforcement learning. The function $\hat{\phi}$ plays a role similar to the value function, and its treatment as a random variable aligns with distributional RL. Learning $\hat{\pi}$ corresponds to policy learning, while $\hat{\Gamma}$ is sometimes referred to as behavior prediction. It is important to note that an agent may adopt more sophisticated strategies by learning additional, related parameters. For example, this may include encoding raw observations, establishing communication protocols, modeling opponents’ incentives, or tuning exploration parameters, such as the best expected rewards in the two-player game example in Section 2.1.
@@
For notation, we didn’t keep track of learning index $n$, but both $\tilde{\mathbb{P}}$ and $\hat{F}$ (and consequently $Q$) are evolving as learning progresses. Note that $Q$ satisfies a Bellman type relation
@@
assigns a scalar value for $(x,a)$, which is better suited in our framework as it does not introduce simplifications. We also note that, in simpler settings, it might be more appropriate to learn a single estimate $Q(x,a)$ directly, and avoid all the estimations we have introduced. This is of course a design choice for the player.
We briefly illustrate a method for learning CartPole [21]. The goal is not to propose a novel learning method, but to further emphasize the connections by learning $\hat{\phi}$ under a fixed horizon $T=8$. Let us revisit the basics of the CartPole problem. The state space is $\mathbb{S}=\mathbb{R}^{4}$ representing position, velocity, angle, and angular velocity. The action space is $\mathbb{A}=\{0,1\}$, where the actions represent applying force to the left or right of the cart. The player gains a reward at each time step if the pole is vertical within some bounds.
The player has a memory for recording observations, including states, actions, and episode-level evaluations. After each episode, a performance metric is computed as the average of the agent’s relative and absolute performance.
@@
To ensure that the value networks are properly trained, we consider the set of all controls $\mathbb{A}^{T}=\{0,1\}^{T}$ and index them by $\{\alpha^{k}\}_{k=1}^{2^{T}}$. [^12] We note that it is considerably easier learning controls in this simple setting, since $\mathbb{A}^{T}$ is finite. Once we include policy learning, the player performs quite well even without properly trained value networks, and hence omitted. Then, we let $\mathfrak{L}_{\pi}:\mathscr{E}\to(\hat{\Omega}\times\mathbb{S}\to{\cal P}(\mathbb{A}^{T}))$ as
@@
Lastly, we will briefly examine the simplest case of Bandit Games and demonstrate how it can be embedded into our framework in two different ways. The first approach adopts a single-player perspective, where the arms are represented by the state $X$. The second approach treats arms as players who take actions but do not engage in learning. Both versions will help to explore our framework further, and may generalize in different directions.
@@
In the action perspective, let players $1,2,\dots,k$ represent arms in the game. Each arm has a fixed horizon $\hat{T}=1$, and transition probabilities or estimating others’ actions are not relevant for the arms. We characterize the reward of the main agent as the action taken by the chosen arm. For the rewards of the arms, one can easily define $\hat{F}^{i}(\hat{\omega},a)$ on a probability space $(\hat{\Omega},\hat{\mathbb{P}})$ such that the induced distributions over actions correspond to the targeted reward distribution $\mu^{i}$.